# An information-matching approach to optimal experimental design and active learning

### v1

@@ Abstract
A model’s predictive performance depends strongly on the quality and quantity of data available for training. Curating comprehensive datasets, however, often confronts practical constraints, including instrumentation, available resources, and cost. Optimal experimental design (OED) [1] and active learning (AL) [2] emerge as practical data collection strategies. Intentionally designing maximally informative experiments guarantees that data are most informative relative to the underlying phenomena of interest, minimize costs, and meet operational requirements. These methodologies have broad applications across scientific domains, including sensor placement problems in power systems [3, 4, 5, 6, 7] and underwater acoustics [8, 9, 10, 10, 11], the development of accurate interatomic potentials in materials science [12, 13, 14, 15], and many other scientific fields [16, 17, 18, 19].
Many OED criteria utilize the Fisher information matrix (FIM), whose inverse establishes a lower bound on parameter covariance, known as the Cramér-Rao bound [20, 21, 22]. Common approaches optimize parameter precision through the FIM, such as by minimizing its trace (A-optimality) [23, 24, 25], or maximizing its determinant (D-optimality) [26, 27, 14, 25] or smallest eigenvalues (E-optimality) [28, 27, 29]. However, many applications of predictive models do not require precise parameter estimates per se, but accurate predictions for key quantities of interest (QoIs) [16, 30, 31, 32]. This distinction is well illustrated by sloppy models, where many parameter combinations are practically unidentifiable, yet still yield precise predictions [33, 34, 35]. Sloppy models exhibit a characteristic information spectrum, where FIM eigenvalues are nearly uniformly spaced on a log scale over many orders of magnitude. Many eigenvalues are smaller than the model’s evaluation precision, rendering the above OED criteria sensitive to numerical noise.
@@
We aim to localize two vertically separated sound sources at depths of 8 and 16 m (red speakers) within $\pm$ 2.5 m vertically and $\pm$ 100 m horizontally (Fig. 3). Candidate receivers are arranged in a rectangular grid (small dots), motivated by common practices of using vertical and horizontal line arrays for ocean sound measurements [8, 9, 10, 11]. We use a range-independent normal mode model, called ORCA [41], to simulate the sound propagation in the ocean and compute the transmission loss at 200 Hz for each candidate location illustrated as small dots in Fig. 3 (details in the SM). The model parameters $\boldsymbol{\theta}$ include the source and receiver locations, as well as the parameterization of the ocean environment. The ocean environment is modeled with 75 m-deep water above an ocean floor consisting of a sandy sediment layer on a half-space basement layer.

### v2

@@ Abstract
A model’s predictive performance depends strongly on the quality and quantity of data available for training. Curating comprehensive datasets, however, often confronts practical constraints, including instrumentation, available resources, and cost. Optimal experimental design (OED) [1] and active learning (AL) [2] emerge as practical data collection strategies. Intentionally designing maximally informative experiments guarantees that data are most informative relative to the underlying phenomena of interest, minimize costs, and meet operational requirements. These methodologies have broad applications across scientific domains, including sensor placement problems in power systems [3, 4, 5, 6, 7, 8] and underwater acoustics [9, 10, 11, 11, 12, 13], the development of accurate interatomic potentials in materials science [14, 15, 16, 17], and many other scientific fields [18, 19, 20, 21].
Many OED criteria utilize the Fisher information matrix (FIM), whose inverse establishes a lower bound on parameter covariance, known as the Cramér-Rao bound [22, 23, 24]. Common approaches optimize parameter precision through the FIM, such as by minimizing its trace (A-optimality) [25, 26, 27], or maximizing its determinant (D-optimality) [28, 29, 16, 27] or smallest eigenvalues (E-optimality) [30, 29, 31]. However, many applications of predictive models do not require precise parameter estimates per se, but accurate predictions for key quantities of interest (QoIs) [18, 32, 33, 34, 35]. This distinction is well illustrated by sloppy models, where many parameter combinations are practically unidentifiable, yet still yield precise predictions [36, 37, 38]. Sloppy models exhibit a characteristic information spectrum, where FIM eigenvalues are nearly uniformly spaced on a log scale over many orders of magnitude. Many eigenvalues are smaller than the model’s evaluation precision, rendering the above OED criteria sensitive to numerical noise.
@@
We aim to localize two vertically separated sound sources at depths of 8 and 16 m (red speakers) within $\pm$ 2.5 m vertically and $\pm$ 100 m horizontally (Fig. 3). Candidate receivers are arranged in a rectangular grid (small dots), motivated by common practices of using vertical and horizontal line arrays for ocean sound measurements [9, 10, 11, 12, 44]. We use a range-independent normal mode model, called ORCA [45], to simulate the sound propagation in the ocean and compute the transmission loss at 200 Hz for each candidate location illustrated as small dots in Fig. 3 (details in the SM). The model parameters $\boldsymbol{\theta}$ include the source and receiver locations, as well as the parameterization of the ocean environment. The ocean environment is modeled with 75 m-deep water above an ocean floor consisting of a sandy sediment layer on a half-space basement layer.